Find the average surface area of a spherical balloon when the radius changes from r = 4 to r = 4.05 cm. The surface area formula for a sphere with radius r is A = 4πr^2 . Round to two decimal places.
Answer #1

solution sunbare area of a spherical ballon о 4- чт 2 when re = yem #, (ул) (2) сут when 1 – 9:09 17 - (5.61 л 4, = (ул) (Ч-o